OpenStudy (anonymous):

which of the following are roots of the following polynomial f(x)=x3-5x2-13x-7

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

math teachers usually make this easy 1 does not work because \[f(1)=1-5-13-7\neq 0\]

OpenStudy (misty1212):

the next good guess is to try \[f(-1)=-1+5+13-7\]

OpenStudy (misty1212):

oops mistake there \[f(-1)=-1-5+13-7\]

OpenStudy (misty1212):

got it on the second try !!

OpenStudy (mathmale):

I see you are trying to answer Dude's question with synthetic division, but have you checked yet whether or not he has studied that before? You might need to explain some preliminaries first.

OpenStudy (mathmale):

Your goal should be to help DUDE get his own answer, not to provide him with the answer.

OpenStudy (misty1212):

?? no synthetic division was harmed in computing the number just plug in and see what you get

OpenStudy (misty1212):

but thanks for your input ...

OpenStudy (mathmale):

Sorry to spill cold water on your fun, but your responses are unacceptable. I am a Moderator and as such it's my job to see that users adhere to OpenStudy rules. Stop now, or risk suspension.

OpenStudy (misty1212):

not sure exactly what the "goal" is since @dude1152 is off line

OpenStudy (misty1212):

guess it must have been that "synthetic division" that was "unacceptable" lol